# Survey on Question Answering over Visually Rich Documents: Methods, Challenges, and Trends

### v1

@@ 3.1 Architectural changes to ViT
A number of approaches leverage CNN architectures, which capture local information more efficiently than ViTs due to their intrinsic design based on convolutions, exploiting equivariant biases inherent to images. Dhouib et al. (2023) proposes a sequential architecture combining CNN and ViT components, where ConvNext blocks are used to extract local features, and their output is fed into a ViT for modeling global dependencies.
@@
Due to the complexity of combining two networks, a CNN and a ViT, without losing information, other approaches Kim et al. (2021); Blecher et al. (2023) draw inspiration from the local window mechanism of CNNs and incorporate it into ViTs, enabling them to process numerous patches effectively. These approaches restrict attention to a local window of patches with a Swin Transformers Liu et al. (2021a), which applies self-attention within local windows, shifting these windows across layers to efficiently integrate cross-window information. However, Swin ViTs progressively reduce the resolution of the tokens through token merging steps, which decrease the number of tokens. DocPedia Feng et al. (2024) removes this downsampling step, enabling it to keep the full token resolution throughout the processing pipeline by leveraging the frequency domain rather than spatially merging patches as done in Swin. Indeed, decomposing an image into the frequency domain, such as using the Discrete Cosine Transform (DCT) Liu et al. (2022a), allows to process larger patches without loosing important high resolution information.
@@
More recent approaches avoid introducing a local bias to ViTs, instead focusing on removing redundant information from ViT patches, as documents often contain a significant amount of redundancy, such as borders, whitespace or decorations. These methods either use attention scores from the self-attention mechanism to prune or merge tokens (e.g., Zhang et al. (2024b); Liu et al. (2023a); Zhao et al. (2023)) or employ unsupervised techniques like Dual-Center K-Means Clustering Zhang et al. (2024a) to compare tokens. TinyChart Zhang et al. (2024b) merges similar tokens at every ViT layer using methods like average pooling, while DocKylin Zhang et al. (2024a) employs similarity-weighted summation based on token cosine similarity ensuring that each token contributes proportionally to its relevance. HRVDA not specified (2024) uses probabilities from a content detector to guide merging, using a shallow segmentation network that focuses on pixel-level information, assigning probabilities to pixels and mapping them to corresponding patches to filter out low-relevance areas (e.g. whitespace or decorations) while preserving meaningful regions (e.g. text or tables). DocKylin Zhang et al. (2024a) goes further in the pixel-level direction of HRVDA not specified (2024), before the image tokenization, to identify redundant information: A preprocessing module (Adaptive Pixel Slimming) modifies the image itself, removing whitespace by detecting low-gradient areas (smooth regions lacking significant content). The identified redundant rows and columns are then cropped, resulting in a more compact and efficient document representation.
These approaches, which modify ViTs by adding local biases (e.g., via CNNs in hybrids or Swin ViTs) or using patch merging/selection, often result in information loss due to merging, pruning, or restricted representations. Recent works have explored pipelines with no modification of the ViT architecture, leveraging pretrained ViTs to process high-resolution images cut into slices, each ViT handling a specific portion of the image.
@@ 3.2 Several ViTs to process partitioned image
The idea of these methods is to slice a high-resolution image into subimages, each processed in parallel by an independent ViT. The resulting representations are combined for a unified document representation.
The way the original image is sliced into subimages is crucial to prevent information loss. Padding preserves the aspect ratio and prevents deformation Li et al. (2024c). Some approaches predict the optimal way to cut the original image, with pre-defined grid matching Ye et al. (2023a) and a score function predicting the best partition Xu et al. (2024), resulting in a varying amount of crop. Whatever the method, models need to maintain the continuity between the different subimages representations.
A first way to do so is through a transformer layer, which helps the model correlate local images to alleviate salient information loss due to cropping Ye et al. (2023a). Liu et al. (2024) introduces a Shifted Window Attention mechanism, enabling sliding window-based attention across subimage representations.
A more efficient approach to maintain continuity between subimages is to leverage a low-resolution document representation to guide the integration of subimages. Through a cross-attention layer, TokenPacker Li et al. (2024c), and later mPLUGDocOwl2 Zhang et al. (2024d), integrate the high-resolution representation of regions into the low-resolution representations using cross-attention, thus interpolating these low-resolution representations with its multi-level region cues treated as reference keys and values to inject their finer information to global image view.
To conclude on vision-only approaches, we think that slicing approaches using local information from cropped image regions to complement a low-resolution global view are promising, enabling compact and efficient representations with significantly fewer tokens while maintaining essential layout and semantic details Zhang et al. (2024d). We also think that techniques like adaptive pixel slimming Zhang et al. (2024a) are promising for VrDU, leveraging the inherent structure of documents, containing whitespace or decorations, to preserve all its content.
@@ 4 Encoding multi-pages documents
Multi-page documents are common in VrDU and are challenging for transformer-based architectures due to their length. Multi-page documents vary in lengths (e.g., 20 pages in SlideVQA Tanaka et al. (2023)), information density (e.g., 21214 in MMLongBench-Doc Ma et al. (2024c)), and cross-page informations (e.g. 2.1% in DUDE Landeghem et al. (2023)). To encode multi-page documents, recent approaches use retrieval-augmented generation (RAG) techniques Lewis et al. (2021) (Section 4.1). Other methods represent the document page by page (Section 4.2), while some rely on long-sequence processing techniques for page interactions (Section 4.3).
@@ 4.1 Retrieval Approach to multi-page
The retrieval approach to multi-page documents focuses on supplying to the VrDU decoder only the representation of pages with relevant information. Several levels can be used to identify the "oracle" element from the document: the retriever can either predict the entire oracle page Naidu et al. (2024); Faysse et al. (2024); Ma et al. (2024b); Cho et al. (2024) or focus on specific regions within the page, such as paragraphs or images that contain the relevant information for answering the question Xie et al. (2024b).
These approaches inherently limit either the interaction between pages or the interaction between modalities, which does not allow cross-page analysis Ma et al. (2024c). Other possibilities, described in the next sections, can be used for multi-modal encoders, but they are however the only viable option to enable vision-only models to handle multi-page inputs (see Table 5), since vision-only models encode a page with many tokens.
4.2 Query-biased approaches
HiVT5 Tito et al. (2023a), and later InstructDr Tanaka et al. (2024), encode each page of the document separately, with a specific learnable token added at the start of each page. HiVT5 Tito et al. (2023a) uses the specialized [PAGE] tokens to guide the encoder in summarizing each document page based on the given question, by processing separately each page with the question, encoding all the relevant information for the next processing step into the [PAGE] token. These [PAGE] tokens representations are then concatenated and passed to the decoder to generate the final answer. To our knowledge, the only vision-only model designed for multi-page input is mPLUG-DocOwl2 Zhang et al. (2024d), which compresses each page representation into 324 tokens and adds a page token for each page. In vision-only approaches, the token length of high-resolution images (i.e., document pages) is typically too large for LLMs to handle multi-page joint understanding, necessitating extreme compression of each page representation and thus degrading performance Zhang et al. (2024d).
However, query-biased approaches only allow limited cross-page reasoning, as the long sequence and diluted information across pages make it challenging to capture specific inter-page relationships Ma et al. (2024c), the page token being not leveraged effectively.
@@ 4.3 Efficient encoding of multi-pages
Another sparse attention approach is implemented by Arctic-Tilt Borchmann et al. (2024), employing a blockwise attention strategy limiting the attention to a chunk size, allowing to handle up to 500 pages (about 390k tokens, or 780 tokens per page on average). This method limits attention to a smaller, predefined neighborhood, reducing complexity from quadratic to linear while representing cross-page information.
An alternative to sparse attention for efficient multi-page documents processing is to use a recurrent network. RM-T5 Gupta et al. (2023) uses a Recurrent Memory Transformer (RMT) Gupta et al. (2022) to process multi-page documents sequentially, treating each page as part of a sequence. This allows the model to carry information across pages by utilizing hidden states from previous pages. The RMT selectively retains or forgets information, capturing essential details from each page for the next encoder, with all memory cells concatenated for the decoder to generate the final answer. However, the drawbacks of RNNs are inherited, such as the lack of parallelization and the limited possible interaction of two elements (here, pages) distant in the sequence.
Overall, we think that approaches that encode entire documents using sparse attention techniques, either global-local or blockwise, represent the future of the multi-page field, as they show great performance on cross-page reasoning Ma et al. (2024c) over retrieval ones.
5 Injecting VrD features into the LLM
Given their high quality generation abilities, large language models (LLMs) have been leveraged in VrDU (see Table 3). Integrating VrD features into an LLM decoder is not straightforward because it requires adapting the VRDs representation space to into an LLM-compatible representation without losing information, while preserving some computational efficiency.
@@ 5.1 Self-attention based approach
This self-attention approach Laurençon et al. (2024) consists in prepending the VrD representation to the prompt, allowing the model to process both VrD features with the prompt together in its self-attention layers. In such approaches, VrD features are projected to the LLM space via several approaches, and are optionally pooled into a shorter sequence.
Methods range from direct linear projection using a single layer to map VrD tokens to the expected input format of the language model Lee et al. (2023a), which projects tokens with minimal parameters, to convolutional ones Yang et al. (2023), allowing to reduce the dimension of the VrD representation, and to using learnable queries to retrieve the relevant VrD tokensLi et al. (2023b); Bai et al. (2023b).
@@
These approaches are limited as they consider the raw tokens of the textual prompt at the same level, without distinguishing between their respective roles or significance.
@@ 5.2 Cross-attention based approach
In the cross-attention-based approach, VrD hidden states encoded by the VrDU encoder are used to condition the frozen LLM using freshly initialized cross-attention layers which are interleaved between the pretrained LLM layers Laurençon et al. (2024). Unlike self-attention, cross-attention approach enables a separate consideration of prompt and VrD tokens. Flamingo Alayrac et al. (2022) pioneered this approach with its Perceiver Resampler, which has since been adopted in various VrDU models (see Table 3).
@@
An advantage of using cross-attention is that it allows to process longer sequences from the encoder, and thus to use only high-resolution representations. For instance, the CogAgent Hong et al. (2023) model employs a high-resolution encoder connected to the decoder through a cross-attention layer: The high-resolution hidden-states condition the LLM.
@@
In other words, cross-attention approaches for integrating visual document (VrD) features into LLMs enable the query/prompt tokens to explicitly interact with VrD features, effectively leveraging the LLM’s capabilities. However, these methods require the introduction of many new parameters, as cross-attention layers are interleaved with the LLM’s architecture, significantly increasing the overall model size Laurençon et al. (2024). We describe this training process in the next section.
6 Training VrDs
VrDU require pre-training for modalities alignment in T+L+V approaches (Section 2), for adapting vision-encoders to text-centric tasks in V-only approaches (Section 3). When using LLMs, additional training is also needed for the interface between the encoder and the LLM.
Model Pretraining Tasks
DocFormer 2021 MVLM, LTR, TDI
TILT 2021 MVLM
LayoutXLM 2021 MVLM, TIA, TIM
SelfDoc 2021b MRP
Pramanik et al. (2022) MVLM, DTM, DSP
LayoutLMv2 2022 MVLM, TIA, TIM
LayoutLMv3 2022 MLM, MIM, WPA
GraphDoc 2022 MSM
ERNIE-Layout 2022 ROP, RRP, MVLM, TIA
UDOP 2023 MVLM, JTL, VTR, MIR
LAMPreT 2021 BOP, MBP, IFP
DocFormerv2 2023 TTL, TTG
GRAM 2024 None
HiVT5 2022b MVLM
DocTr 2023 MDM
Arctic-TILT 2024 MVLM
LayoutLLM 2024 MLM, MIM, WPA
Table 4: Pretraining tasks for modalities alignment.
6.1 Pretraining for modalities alignment
Many techniques have been proposed to align multiple modalities, leveraging the fact that the information is somehow redundant (when text and image is present), allowing to align the multiple modalities. The different cross-modal pretraining tasks in VrDU models aim to align textual, positional, and visual modalities, are summarized in Table 4.
Inspired by the masked language modeling (MLM) training approach in NLP, the Masked Visual Language Modeling (MVLM) was introduced by LayoutLM Xu et al. (2019), which masks text but not the layout information. Another possibility is to mask bounding boxes while keeping the text visible, as done in StructuralLM Li et al. (2021a). Joint text-layout (JTL) prediction Tang et al. (2023) and Area-Masked Language Modeling (AMLM) Hong et al. (2021) mask both text and layout. Masked Detection Modeling (MDM) Liao et al. (2023), Text-Layout Pretraining (TLP) Mao et al. (2024) and Layout Prediction (LP) Tang et al. (2023) predict masked text and layout using visual features.
Inspired by Masked Autoencoders (MAE) He et al. (2021a), which mask large portions of an input image and train the model to reconstruct the missing parts, several approaches have adapted this concept to capture finer visual and contextual details in document understanding. Masked Image Modeling (MIM) Huang et al. (2022) leverages discrete Variational Autoencoders (dVAE) to transform image patches into tokens, allowing token-level comparison to reconstruct masked regions. Building on this, Masked Image Reconstruction (MIR) Tang et al. (2023); Li et al. (2023e) directly predicts RGB pixel values for masked document regions. The Learn to Reconstruct (LTR) Appalaraju et al. (2021) task does the same for the whole image. Masked Region Modeling (MRM) Chen et al. (2020) incorporates both text and visible regions to infer the masked areas, while Image Fitting (IMG-FIT) Wu et al. (2021) shifts the focus to selecting the correct image from multiple options, providing a discriminative task.
Contrastive strategies train models to match or distinguish modality pairs by adjusting their similarity in a shared space. DoCo Li et al. (2024d) aligns text tokens and bounding boxes with visual features to connect words to their spatial context. LayoutLMv2 Xu et al. (2022) introduced Text-Image Alignment (TIA) to align text groups with image regions, using binary cross-entropy to predict if regions contain the text. LayoutLMv3 Huang et al. (2022) improved this with Word-Patch Alignment (WPA), aligning individual text tokens with image patches from a ViT. Region Alignment (WRA) Chen et al. (2020) uses Optimal Transport (OT) to align text tokens and image regions by finding a transport plan that matches word and visual region distributions, minimizing the alignment cost measured by cosine distance.
Captioning Strategies (IC) can be used to align text and image. Text Describes Image (TDI) task, used in DocFormer Appalaraju et al. (2021), aims at verifying if text matches a document image. Pramanik et al. (2022) later proposed Document Topic Modeling (DTM), using semantic contexts as soft labels to learn visual features and represent documents as topic distributions.
Finally, as layout does not translate to unambiguous reading order of the document, pretraining tasks linked to this aspect include Reading Order Prediction (ROP) Peng et al. (2022), training the model to recognize correct order between blocks, or Block-Ordering Predictions (B-ORD) Wu et al. (2021), which swaps two content blocks for the model to predict if the order is correct. Token-to-Grid (TTG) Appalaraju et al. (2023) divides the document into a grid and predicts which token belongs to which cell. Pramanik et al. (2022) extend reading-order tasks to multi-pages with Document Shuffle Prediction (DSP) task, training the model to detect shuffled pages in multi-page documents.
6.2 Pretraining Vision-only models
@@
Hu et al. (2024a) highlight that vision-only models must be pre-trained on document parsing tasks to be capable of understanding the document content. To do so, it is possible to exploit the fact that documents are often generated from a symbolic source document (e.g. LaTeX, HTML), so the pretraining task aims at converting images into a structured text format: Webpage screenshots into HTML Lee et al. (2023a), PDF pages into LaTeX Wei et al. (2023) or Markdown Blecher et al. (2023), and table/chart into extended Markdown formats Hu et al. (2024a), CSV or JSON Kim et al. (2021). Hu et al. (2024a) implements a multi-format hierarchical reconstruction task named Unified Structured Learning, using multiple publicly available datasets and constructing structure-aware text sequences or text-position pairs as the targets.
We note that this pre-training task does not necessarily require an LLM decoder; to conserve resources, a lightweight decoder could be used during this phase Wei et al. (2024). However, most approaches leverage this pre-training task to align VrD features with the LLM decoder, training not only the visual encoder but also the projector Hu et al. (2024a).
Models Doc Info Chart DUDE MPDoc MMLong
VQA VQA QA VQA Bench
Multimodal (text, layout, visual) VrDU (Sec 2)
TILT 2021 87.1
@@
DocFormerv2 2023 87.8 48.8 50.8² 76.8²
HiVT5 2023a 23.0 62.0
LayoutLLM 2024 86.9
DocLayLLM 2024 78.4 40.9
GRAM 2024 86.0 53.4 80.3
Arctic-TILT 2024 90.2 57.0 58.1 81.2 25.8
Vision-only VrDU models (Sec 3)
DONUT 2021 72.1 11.6 41.8
@@
Pix2Struct 2023a 76.6 40.0 58.6 62.0*
SeRum 2023b 77.9
LLaVAR 2023 6.73 12.3 13.5
Unidoc 2023 7.70 14.7 10.88
mPLUGDoc 2023b 62.2 38.2 57.4
QwenVL 2023a 65.1 35.4 65.7 84.4* 6.1*
UREADER 2023b 65.4 42.2 59.3
Monkey 2023 66.5 36.1 65.1 6.2*
TextSquare 2024 84.3 51.5 79.6
TextMonkey 2024 73.0 28.6 66.9
TextHawk 2024 76.4 50.6 66.6
DocPedia 2023a 47.8 15.2 46.9
CogAgent 2023b 81.6 44.5 68.4
Vary 2023a 76.3 66.1
mPLUGDoc1.5 2024a 82.2 50.7 70.2 6.9*
TinyChart 2024b 83.6
ILMXC24KHD 2024b 90.0 68.6 81.0 56.1* 76.9* 10.3*
Idefics2 2024 74.0 56.0* 7.0*
TokenPacker 2024c 70.0
mPLUGDoc2 2024d 80.7 46.4 70.0 46.8 69.4
Closed Models
GPT-4V 88.4 75.1 78.5
GPT-4o 92.8 85.5 54.0 67.0 42.8

### v2

@@ 3.1 Architectural changes to ViT
A number of approaches leverage CNN architectures, which capture local information more efficiently than ViTs due to their intrinsic design based on convolutions, exploiting locality bias in images. Dhouib et al. (2023) proposes a sequential architecture combining CNN and ViT components, where ConvNext blocks are used to extract local features, and their output is fed into a ViT for modeling global dependencies.
@@
Due to the complexity of combining two networks without losing information, other approaches Kim et al. (2022); Blecher et al. (2023) draw inspiration from the local window mechanism of CNNs and incorporate it into ViTs, enabling them to process numerous patches effectively. These approaches restrict attention to a local window of patches with a Swin Transformers Liu et al. (2021), which applies self-attention within local windows, shifting these windows across layers to efficiently integrate cross-window information. However, Swin ViTs progressively reduce the resolution of the tokens through token merging steps, which decrease the number of tokens. DocPedia Feng et al. (2024) removes this downsampling step, keeping the full token resolution throughout the processing pipeline by leveraging the frequency domain rather than spatially merging patches as done in Swin. More precisely, they represent an image in the frequency domain, using the Discrete Cosine Transform Liu et al. (2022a), allowing to process larger patches without loosing important high resolution information. However, restricting the attention to local windows, even if shifted, introduces a locality bias to ViTs, similar to CNNs.
@@
More recent approaches avoid introducing a locality bias to ViTs, instead focusing on removing redundant information from ViT patches, as documents often contain a significant amount of redundancies, such as borders, whitespace or decorations. These methods either use attention scores from the self-attention mechanism to prune or merge tokens (e.g., Zhang et al. (2024b); Shang et al. (2024); Chen et al. (2024)) or employ unsupervised techniques like Dual-Center K-Means Clustering Zhang et al. (2024a) to select tokens. TinyChart Zhang et al. (2024b) combines similar tokens after each ViT layer using methods like average pooling, while DocKylin Zhang et al. (2024a) employs similarity-weighted summation based on token cosine similarity ensuring that each token contributes proportionally to its relevance. Other approaches Liu et al. (2024a) use a content detection module to filter out low-relevance areas (e.g., whitespace) and preserve meaningful regions (e.g., text or tables) by assigning probabilities to pixels and mapping them to patches.
@@ 3.2 Several ViTs to process partitioned image
Recent works have explored pipelines leveraging already pretrained ViTs to process high-resolution images cut into slices. Each ViT handles a specific portion of the image, and the resulting representations are combined (sequence of "image tokens") as the unified document representation.
The way the original image is sliced into subimages is crucial to prevent information loss. Padding preserves the aspect ratio and prevents deformation Li et al. (2024b). Some approaches predict the optimal way to cut the original image, with pre-defined grid matching Ye et al. (2023b) and a score function predicting the best partition Xu et al. (2024b), resulting in a varying amount of crop. Whatever the method, models need to maintain the continuity between the different subimages representations.
A simple way to do so is through a 2D crop position encoding, which allows interaction between local images Ye et al. (2023b). However, this approach lacks information continuity between cropped images. To alleviate salient information loss due to cropping, Liu et al. (2024b) introduces a Shifted Window Attention mechanism, enabling sliding window-based attention across subimage representations.
A more efficient approach to maintain continuity between subimages is to leverage a low-resolution document representation to guide the integration of subimages. Through a cross-attention layer, TokenPacker Li et al. (2024b), and later mPLUGDocOwl2 Hu et al. (2024b), integrate the high-resolution representation of regions into the low-resolution representations using cross-attention, thus interpolating these low-resolution representations with its multi-level region cues treated as reference keys and values to inject their finer information to global image view.
To conclude on vision-only approaches, we think that slicing approaches using local information from cropped image regions to complement a low-resolution global view are promising, enabling compact and efficient representations with significantly fewer tokens while maintaining essential layout and semantic details Hu et al. (2024b). However, while this type of approach reduces computational cost for single-page processing, it is not sufficient to handle multi-page Hu et al. (2024b).
@@ 4 Encoding multi-pages documents
The principal challenge in VrDU is to handle multi-pages documents. Multi-page documents vary in length (e.g., 20 pages in SlideVQA Tanaka et al. (2023)), amount of tokens per document (e.g., 21214 tokens per document in MMLongBench-Doc Ma et al. (2024c)), and cross-page information, i.e. questions requiring information from several pages of the document (e.g. 2.1% in DUDE Landeghem et al. (2023)). To encode multi-page documents, recent approaches use retrieval-augmented generation (RAG) techniques Lewis et al. (2021) (Section 4.1). Other methods represent the document page by page (Section 4.2), enhanced with inter-page interactions inherited from long-sequence processing techniques (Section 4.3).
@@ 4.1 Retrieval Approach to multi-page
The retrieval approach to multi-page documents focuses on supplying to the VrDU decoder only the representation of pages with relevant information. Several levels can be used to identify the relevant element from the document: the retriever can either predict the entire relevant page Naidu et al. (2024); Faysse et al. (2024); Ma et al. (2024b); Cho et al. (2024) or focus on specific regions within the page, such as paragraphs or images containing the elements to answer the question Xie et al. (2024).
These approaches inherently limit either the interaction between pages or the interaction between modalities, which does not allow cross-page analysis Ma et al. (2024c), not mentioning that they highly depend on the performance of the retriever.
4.2 Query-based approaches
HiVT5 Tito et al. (2023), and later InstructDr Tanaka et al. (2024), encode each page of the document separately, with a specific learnable token added at the start of each page. HiVT5 Tito et al. (2023) uses the specialized [PAGE] tokens to guide the encoder in summarizing each document page based on the given question, by processing separately each page with the question, encoding all the relevant information for the next processing step into the [PAGE] token. These [PAGE] tokens representations are then concatenated and passed to the decoder to generate the final answer. To our knowledge, the only vision-only model designed for multi-page input is mPLUG-DocOwl2 Hu et al. (2024b), which compresses each page representation into 324 tokens and adds a page token for each page. In vision-only approaches, the token length of high-resolution images (i.e., document pages) is typically too large for LLMs to handle multi-page joint understanding, necessitating extreme compression of each page representation and thus degrading performance Hu et al. (2024b).
However, query-based approaches only allow limited cross-page reasoning, as the long sequence and diluted information across pages make it challenging to capture specific inter-page relationships Ma et al. (2024c), the page token being not leveraged effectively.
@@ 4.3 Efficient encoding of multi-pages
Another sparse attention approach is implemented by Arctic-Tilt Borchmann et al. (2024), employing a blockwise attention strategy limiting the attention to a chunk size, allowing to handle up to 500 pages (about 390k tokens, with 780 tokens per page on average). This method limits attention to a smaller, predefined neighborhood ($\approx$2 pages), reducing complexity from quadratic to linear while representing cross-page information.
An alternative to sparse attention for efficient multi-page documents processing is to use a recurrent network. RM-T5 Dong et al. (2024a) uses a Recurrent Memory Transformer (RMT) Gupta et al. (2022) to process multi-page documents sequentially, treating each page as part of a sequence. This allows the model to carry information across pages by utilizing hidden states from previous pages. The RMT selectively retains or forgets information, capturing essential details from each page for the next encoder, with all memory cells concatenated for the decoder to generate the final answer. However, the drawbacks of RNNs are inherited, such as the lack of parallelization and the limited possible interaction of two elements (here, pages) distant in the sequence.
Overall, our view is that approaches that encode entire documents using sparse attention techniques, either global-local or blockwise, represent the future of the multi-page field, as they show great performance on cross-page reasoning Ma et al. (2024c) over retrieval ones.
5 Injecting visual features into the LLM
In both approaches for encoding the VrD (structured encoding in Section 2 versus vision-only encoding in Section 3), the representation of the document contains visual features. Integrating visual features into an LLM decoder is not straightforward because it requires adapting the visual representation space into an LLM-compatible representation without losing information, while preserving some computational efficiency. We detail here how this integration is done by current VrDU approaches, and what the future directions for visual features integration into LLMs are.
@@ 5.1 Self-attention based approach
This self-attention approach Laurençon et al. (2024) consists in prepending the visual representation to the prompt, allowing the model to process both visual features with the prompt together in its self-attention layers. In such approaches, visual features are projected into the LLM space via several approaches, and are optionally pooled into a shorter sequence.
Those methods vary in complexity, ranging from direct linear projection using a single layer to map visual tokens to the expected input format of the language model Lee et al. (2023), which minimizes the number of parameters; convolutional approaches Cha et al. (2024), which reduce the dimensionality of the visual representation; to using learnable queries Li et al. (2023a); Bai et al. (2023), used to retrieve relevant visual tokens.
@@
These approaches are limited, considering raw tokens of the textual prompt and visual tokens from the document at the same level, without distinguishing between their respective roles or significance.
@@ 5.2 Cross-attention based approach
In the cross-attention-based approach, visual hidden states encoded by the visual encoder are used to condition a frozen LLM using freshly initialized cross-attention layers which are interleaved between the pretrained LLM layers Laurençon et al. (2024). Unlike self-attention, cross-attention approach enables a separate consideration of prompt and visual document tokens. Flamingo Alayrac et al. (2022) pioneered this approach with its Perceiver Resampler, which has since been adopted in various VrDU models (see Table 2).
@@
An advantage of using cross-attention is that it allows to process longer sequences from the encoder, and thus to use only high-resolution representations. For instance, CogAgent Hong et al. (2024) employs a high-resolution encoder connected to the decoder through a cross-attention layer, while using self-attention with a low resolution version of the image.
@@
In other words, cross-attention approaches for integrating visual features into LLMs enable the query/prompt tokens to explicitly interact with visual features, effectively leveraging the LLM’s capabilities.
However, these methods require the introduction of many new parameters, as cross-attention layers are interleaved with the LLM’s architecture, significantly increasing the overall model size Laurençon et al. (2024).
5.3 Pretraining for visual features insertion
@@
Hu et al. (2024a) highlight that, to integrate visual features into an LLM, VrDU models must be pretrained on document parsing tasks. Lee et al. (2023); Wei et al. (2023); Blecher et al. (2023); Hu et al. (2024a); Kim et al. (2022) exploit the fact that documents are often generated from a symbolic source document (e.g. HTML, latex, Markdown, extended Markdown format for table and charts or CSV/JSON) to convert document page screenshot into structured text for pretraining. Hu et al. (2024a) implements a multi-format reconstruction task named Unified Structured Learning.
Models Doc Info DUDE MPDoc
VQA VQA VQA
T+L+V models (Section 2)
@@
DocFormerv2 2023 87.8 48.8 50.8² 76.8²
HiVT5 2023 23.0 62.0
GRAM 2024 86.0 53.4 80.3
LayoutLLM 2024 86.9
DocLayLLM 2024 78.4 40.9
TILT 2021 87.1
@@
Arctic-TILT 2024 90.2 57.0 58.1 81.2
Vision-only models (Section 3)
DONUT 2022 72.1 11.6
@@
Pix2Struct 2023 76.6 40.0 62.0*
SeRum 2023 77.9
Kosmos2.5 2024 81.1 41.3
LLaVAR 2024d 6.73 12.3
Unidoc 2023 7.70 14.7
DocPedia 2024 47.8 15.2
CogAgent 2024 81.6 44.5
Vary 2023 76.3
mPLUGDoc 2023a 62.2 38.2
QwenVL 2023 65.1 35.4 84.4*
UREADER 2023b 65.4 42.2
Monkey 2024d 66.5 36.1
TextSquare 2024 84.3 51.5
TextMonkey 2024b 73.0 28.6
mPLUGDoc1.5 2024a 82.2 50.7
ILMXC24KHD 2024b 90.0 68.6 56.1* 76.9*
Idefics2 2024 74.0 56.0*
TextHawk 2024 76.4 50.6
TokenPacker 2024b 70.0
mPLUGDoc2 2024b 80.7 46.4 46.8 69.4
HRVDA 2024a 72.1 43.5
DocKylin 2024a 77.3 46.6
Commercial Models
GPT-4V 88.4 75.1
GPT-4o 92.8 54.0 67.0